# Tropicalization of curve arrangement complements and arroids

### v2

@@ 7. Maximal subvarieties
[…]
@@
Let $\mathcal{B}$ be a simple arrangement of curves, with all intersection points being real, and with the corresponding arroid fan $\Sigma_{\mathbf{A}_{\mathcal{B}}}$ being a tropical homology manifold. By Proposition 6.1, the arrangement complement $X_{\mathcal{B}}$ is wunderschön, and so by [2, Theorem 6.1], it is cohomologically tropical. By [2, Theorem 6.2], we have isomorphisms $H^{k}(X_{\mathcal{B}}(\mathbb{C}))\cong H^{k,0}(\Sigma_{\mathbf{A}_{\mathcal{B}}})$, and in particular $b_{k}(X_{\mathcal{B}}(\mathbb{C});\mathbb{Z}/2\mathbb{Z})=\dim\;H^{k,0}(\Sigma_{\mathbf{A}_{\mathcal{B}}})$, for all $k$ since the homology is torsion-free by Lemma 7.1.
@@
For any choice of curve $C\in\mathcal{B}$ and corresponding $i\in\mathcal{A}$, the reduced star fan of the ray $\rho_{i}$ is given by $\Sigma_{\mathbf{A}/i}$ by Proposition 4.13, and is a tropical homology manifold by assumption. It follows from Proposition 6.1 that the punctured curve $C^{*}$ tropicalizing to $\Sigma_{\mathbf{A}/i}$ is wunderschön, and so by [2, Theorem 6.2], $\dim\,H^{k}(C^{*})=\dim\,H^{k,0}(\Sigma_{\mathbf{A}/i})$ for all $k$. Moreover, $C^{*}(\mathbb{C})$ is homotopic to a wedge of circles, so that its homology is torsion-free, hence $b_{k}(C^{*}(\mathbb{C}))=b_{k}(C^{*}(\mathbb{C});\mathbb{Z}/2\mathbb{Z})=\dim\;H^{k,0}(\Sigma_{\mathbf{A}/i})$ for all $k$.
@@
Similarly, we study the arrangement $\mathcal{B}\setminus\{C\}$, corresponding to the contracted arroid $\mathbf{A}_{\mathcal{B}\setminus i}$. The tropicalization of its complement $X_{\mathcal{B}\setminus\{C\}}$ is the arroid fan $\Sigma_{\mathbf{A}_{\mathcal{B}\setminus i}}$, which by Lemma 5.5 is a tropical homology manifold. By induction assumption, the complement $X_{\mathcal{B}\setminus\{C\}}$ is maximal, i.e. $b_{\bullet}(X_{\mathcal{B}\setminus\{C\}}(\mathbb{R}))=b_{\bullet}(X_{\mathcal{B}\setminus\{C\}}(\mathbb{C}))$. By Proposition 6.1, $X_{\mathcal{B}\setminus\{C\}}$ is wunderschön, so that by [2, Theorem 6.1] it is cohomologically tropical, and therefore $H^{k,0}(\Sigma_{\mathbf{A}_{\mathcal{B}\setminus i}})\cong H^{k}(X_{\mathcal{B}\setminus\{C\}}(\mathbb{C}))$ for all $k$ by [2, Theorem 6.2]. Therefore $b_{k}(X_{\mathcal{B}\setminus\{C\}}(\mathbb{C}))=b_{k}(X_{\mathcal{B}\setminus\{C\}}(\mathbb{C});\mathbb{Z}/2\mathbb{Z})=\dim\;H^{k,0}(\Sigma_{\mathbf{A}_{\mathcal{B}\setminus i}})$ for all $k$.
@@
By Lemma 7.2, $b_{\bullet}(X_{\mathcal{B}}(\mathbb{R}))=b_{0}(X_{\mathcal{B}}(\mathbb{R}))$ for any simple arrangement with real intersection points. Moreover, each arc of $C^{*}(\mathbb{R})$ increases the number of connected components of $X_{\mathcal{B}\setminus\{C\}}(\mathbb{R})$ by $1$, so counting the number of connected components of $X_{\mathcal{B}}(\mathbb{R})$, we have $b_{0}(X_{\mathcal{B}}(\mathbb{R}))=b_{0}(X_{\mathcal{B}\setminus\{C\}}(\mathbb{R}))+b_{0}(C^{*}(\mathbb{R}))$. Note also that the number of punctures of $S^{1}\cong C(\mathbb{R})$ by real points is $b_{0}(C^{*}(\mathbb{R}))$, so $C^{*}(\mathbb{C})$ is a punctured Riemann sphere, which gives $b_{0}(C^{*}(\mathbb{R}))=b_{\bullet}(C^{*}(\mathbb{C}))$. Summarizing, we have
@@
By Lemma 5.2, we have $\dim\,H^{k-1,0}(\Sigma_{\mathbf{A}/i})+\dim\,H^{k,0}(\Sigma_{\mathbf{A}\setminus i})=\dim\,H^{k,0}(\Sigma_{\mathbf{A}})$ for all $k$, which when combined with the equalities between tropical cohomology and singular cohomology described above, yield $b_{\bullet}(X_{\mathcal{B}\setminus\{C\}}(\mathbb{C}))+b_{\bullet}(C^{*}(\mathbb{C}))=b_{\bullet}(X_{\mathcal{B}}(\mathbb{C}))$. ∎
@@
We illustrate Theorem 7.3 by the example displayed in Figure 6. It is constructed by selecting four points, drawing all lines between them, and then selecting any number of non-singular conics passing through those same four points. To see that these arrangements are tropical homology manifolds, it suffices to note that in each intersection point of a line with a conic, multiple other lines pass through the same point, so that all rays associated to lines of the arroid fan are uniquely balanced by Lemma 5.10, and all intersection points of any pair of conics are connected by lines, so that by Lemma 5.13 all rays associated to conics are uniquely balanced. By Theorem 5.4, the arroid fan is therefore a tropical homology manifold, hence by Theorem 7.3 each such arrangement is maximal.
@@
The first property follows from Lemma 7.1 and Lemma 7.2, the second from Theorem 7.3. The purity of the mixed Hodge structure on rational cohomology is satisfied because $X_{\mathcal{B}}$ is wunderschön by Proposition 6.1, and the integer homology is torsion-free by Lemma 7.1. The last property follows from Theorem 6.2, as $X_{\mathcal{B}}$ is cohomologically tropical since the tropicalization is a tropical homology manifold. ∎
In particular, the family of arrangements described in Example 7.4 gives examples of the varieties satisfying the conditions of [4].

### v3

@@ 7. Maximal subvarieties
$\cdots$$H_{3}(\mathbb{P}_{\mathbb{C}}^{2};\mathbb{Z})$$H_{3}(\mathbb{P}_{\mathbb{C}}^{2},X_{\mathcal{B}}(\mathbb{C});\mathbb{Z})$$H_{2}(X_{\mathcal{B}}(\mathbb{C});\mathbb{Z})$$H_{2}(\mathbb{P}_{\mathbb{C}}^{2};\mathbb{Z})$$\cdots$}}
@@
Let $\mathcal{B}$ be a simple arrangement of curves, with all intersection points being real, and with the corresponding arroid fan $\Sigma_{\mathbf{A}_{\mathcal{B}}}$ being a tropical homology manifold. By Proposition˜6.1, the arrangement complement $X_{\mathcal{B}}$ is wunderschön, and so by [2, Theorem 6.1], it is cohomologically tropical. By [2, Theorem 6.2], we have isomorphisms $H^{k}(X_{\mathcal{B}}(\mathbb{C}))\cong H^{k,0}(\Sigma_{\mathbf{A}_{\mathcal{B}}})$, and in particular $b_{k}(X_{\mathcal{B}}(\mathbb{C});\mathbb{Z}/2\mathbb{Z})=\dim\;H^{k,0}(\Sigma_{\mathbf{A}_{\mathcal{B}}})$, for all $k$ since the homology is torsion-free by Lemma˜7.1.
@@
For any choice of curve $C\in\mathcal{B}$ and corresponding $i\in\mathcal{A}$, the reduced star fan of the ray $\rho_{i}$ is given by $\Sigma_{\mathbf{A}/i}$ by Proposition˜4.13, and is a tropical homology manifold by assumption. It follows from Proposition˜6.1 that the punctured curve $C^{*}$ tropicalizing to $\Sigma_{\mathbf{A}/i}$ is wunderschön, and so by [2, Theorem 6.2], $\dim\,H^{k}(C^{*})=\dim\,H^{k,0}(\Sigma_{\mathbf{A}/i})$ for all $k$. Moreover, $C^{*}(\mathbb{C})$ is homotopic to a wedge of circles, so that its homology is torsion-free, hence $b_{k}(C^{*}(\mathbb{C}))=b_{k}(C^{*}(\mathbb{C});\mathbb{Z}/2\mathbb{Z})=\dim\;H^{k,0}(\Sigma_{\mathbf{A}/i})$ for all $k$.
@@
Similarly, we study the arrangement $\mathcal{B}\setminus\{C\}$, corresponding to the contracted arroid $\mathbf{A}_{\mathcal{B}\setminus i}$. The tropicalization of its complement $X_{\mathcal{B}\setminus\{C\}}$ is the arroid fan $\Sigma_{\mathbf{A}_{\mathcal{B}\setminus i}}$, which by Lemma˜5.5 is a tropical homology manifold. By induction assumption, the complement $X_{\mathcal{B}\setminus\{C\}}$ is maximal, i.e. $b_{\bullet}(X_{\mathcal{B}\setminus\{C\}}(\mathbb{R}))=b_{\bullet}(X_{\mathcal{B}\setminus\{C\}}(\mathbb{C}))$. By Proposition˜6.1, $X_{\mathcal{B}\setminus\{C\}}$ is wunderschön, so that by [2, Theorem 6.1] it is cohomologically tropical, and therefore $H^{k,0}(\Sigma_{\mathbf{A}_{\mathcal{B}\setminus i}})\cong H^{k}(X_{\mathcal{B}\setminus\{C\}}(\mathbb{C}))$ for all $k$ by [2, Theorem 6.2]. Therefore $b_{k}(X_{\mathcal{B}\setminus\{C\}}(\mathbb{C}))=b_{k}(X_{\mathcal{B}\setminus\{C\}}(\mathbb{C});\mathbb{Z}/2\mathbb{Z})=\dim\;H^{k,0}(\Sigma_{\mathbf{A}_{\mathcal{B}\setminus i}})$ for all $k$.
@@
By Lemma˜7.2, $b_{\bullet}(X_{\mathcal{B}}(\mathbb{R}))=b_{0}(X_{\mathcal{B}}(\mathbb{R}))$ for any simple arrangement with real intersection points. Moreover, each arc of $C^{*}(\mathbb{R})$ increases the number of connected components of $X_{\mathcal{B}\setminus\{C\}}(\mathbb{R})$ by $1$, so counting the number of connected components of $X_{\mathcal{B}}(\mathbb{R})$, we have $b_{0}(X_{\mathcal{B}}(\mathbb{R}))=b_{0}(X_{\mathcal{B}\setminus\{C\}}(\mathbb{R}))+b_{0}(C^{*}(\mathbb{R}))$. Note also that the number of punctures of $S^{1}\cong C(\mathbb{R})$ by real points is $b_{0}(C^{*}(\mathbb{R}))$, so $C^{*}(\mathbb{C})$ is a punctured Riemann sphere, which gives $b_{0}(C^{*}(\mathbb{R}))=b_{\bullet}(C^{*}(\mathbb{C}))$. Summarizing, we have
@@
By Lemma˜5.2, we have $\dim\,H^{k-1,0}(\Sigma_{\mathbf{A}/i})+\dim\,H^{k,0}(\Sigma_{\mathbf{A}\setminus i})=\dim\,H^{k,0}(\Sigma_{\mathbf{A}})$ for all $k$, which when combined with the equalities between tropical cohomology and singular cohomology described above, yield $b_{\bullet}(X_{\mathcal{B}\setminus\{C\}}(\mathbb{C}))+b_{\bullet}(C^{*}(\mathbb{C}))=b_{\bullet}(X_{\mathcal{B}}(\mathbb{C}))$. ∎
@@
We illustrate Theorem˜7.3 by the example displayed in Figure˜6. It is constructed by selecting four points, drawing all lines between them, and then selecting any number of non-singular conics passing through those same four points. To see that these arrangements are tropical homology manifolds, it suffices to note that in each intersection point of a line with a conic, multiple other lines pass through the same point, so that all rays associated to lines of the arroid fan are uniquely balanced by Lemma˜5.10, and all intersection points of any pair of conics are connected by lines, so that by Lemma˜5.13 all rays associated to conics are uniquely balanced. By Theorem˜5.4, the arroid fan is therefore a tropical homology manifold, hence by Theorem˜7.3 each such arrangement is maximal.
@@
The first property follows from Lemma˜7.1 and Lemma˜7.2, the second from Theorem˜7.3. The purity of the mixed Hodge structure on rational cohomology is satisfied because $X_{\mathcal{B}}$ is wunderschön by Proposition˜6.1, and the integer homology is torsion-free by Lemma˜7.1. The last property follows from Theorem˜6.2, as $X_{\mathcal{B}}$ is cohomologically tropical since the tropicalization is a tropical homology manifold. ∎
In particular, the family of arrangements described in Example˜7.4 gives examples of the varieties satisfying the conditions of [4].